An investor has $50,000 to invest in three types of bonds: short-term, intermediate-term, and long-term. How much should she invest in each type to satisfy the given conditions?
Short-term bonds pay 4% annually, intermediate-term bonds pay 5%, and long-term bonds pay 6%. The investor wishes to realize a total annual income of 4.92%, with equal amounts invested in short- and intermediate-term bonds.
short-term bonds $
intermediate-term bonds $
long-term bonds $

Answer by richwmiller(17219)

x+y+z=50000
0.04*x+0.05*y+0.06*z=2460
we have two equations in three unknown we need one more equation.
x=y
since x=y
2x+z=50000
0.04*x+0.05*x+0.06*z=2460
0.09*x+0.06*z=2460
2x+z=50000
z=50000-2x
This won't be pretty but it can be solved.
0.09*x+0.06*(50000-2x) =2460
0.09*x+0.06*(50000-2x) =2460
0.09*x+0.06*50000-.06*2x =2460
0.09*x+3000-.12x =2460
-.03x=-540
multiply by -100
3x=54000
x=18000
y=18000
z=14000